Problem #3: Bond concept Presented below are a variety of questions. Select the best answer choice for each multiple-choice question. Indicate your selection by writing the corresponding letter in the space provided at the bottom right of the question. 1. On May 5, 2018, Bergeron Investment Company purchased $10 million face value of 15-year, 7.5% bonds issued by Boychuk Defense Systems at a price of 99.5. a. Bergeron Investment Company purchased the Boychuk bonds at a market yield rate that was: equal to 7.5% greater than 7.5% A. C, less than 7.5% D. cannot be determined from the information provided Your Answer 2 points b. On June 30, 2018, Bergeron Investment Company determined the Boychuk bonds were selling at a price that yielded investors 7.3%. The price of the bond on June 30, 2018 was: A. less than 99.5 B. equal to 100 C. more than 100 D. cannot be determined from the information provided 2 points Your Answer c. Bergeron Investment Company sold the Boychuk bonds at 102.5 on July 31, 2018. Compute the capital gain that Bergeron realized on the investment: A. $30,000,000 B. $3,000,000 C. $300,000 D. cannot be determined from the information provided 2 points Your Answer
Answer #1

Please provide rating....

Answer 1)a.
Correct answer is option B. Greater than 7.5%
Since the bond is issued at discount market yield is higher than 7.5%
Answer 1)b.
Correct answer is option C. More than 100
Since bond yield is less than 7.5% bone will be traded at premium
Answer 1)c.
Correct answer is option C. 300,000
a Number of bond purchased =        100,000
b Purchase price = 99.5
c Sale price = 102.5
d=c-b Profit per bond = 3
e=d*a Total profit =        300,000